### v3

@@ 6 Counterfactual program assignment policies
The number of spots in each program is limited. The capacity of Mentoring is constrained by the availability of mentors, while in Challenges, we capped spots to 300 to accommodate potential operational complexities and logistical issues. Considering the substantial treatment effects of both programs, one might posit that relaxing these capacity constraints will create significant value. Nonetheless, capacity constraints will likely remain a long-term issue for the mentoring program and may persist, albeit to a lesser degree, for Challenges. Consequently, Dare IT will continue being faced with the question of which applicants to accept for each program. With heterogeneous treatment effects, admission policies favoring applicants with high predicted benefits could enhance Dare IT programs’ impact. This section aims to quantify such potential gains.
@@
We carry out three analyses: (i.) we estimate gains from the introduction of a policy that, for given capacity levels in each program, assigns applicants to programs to maximize the sum of predicted treatment effects across the two programs; (ii.) we compare the overall impact of Mentoring when mentors select most promising candidates to a case when they select to maximize expected benefits; ((iii). finally, we evaluate several simple policies based on heuristics.
@@
An assignment policy is a mapping from applicants’ characteristics $X\in\mathcal{X}$ and programs capacity levels $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT. $\mathcal{Q}$ is the grid of possible capacity levels under consideration. Policy estimation and evaluation are carried out separately. The objective of policy estimation is to obtain a mapping from user characteristics and programs’ capacity level to program assignments’. This mapping $\hat{\pi}$ is estimated in a training set. Across all analyses, we obtain training and testing sets by randomly splitting data at the applicant level with equal probability. Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping to the testing set, where for each $i$ in the test set and each $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. Second, for each $Q\in\mathcal{Q}$, we construct the AIPW estimates of the “policy value,” that is, the value of the outcome when treatments are assigned according to $\pi(X_{i},Q)\in\mathcal{P}$. We do that using cross-fitting in the testing sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds (we use 100 repetitions with 10 folds and consider the median outcome across repetitions) [^16] We follow the approach suggested in Chernozhukov et al. (2017).. Throughout this section, we use Cox proportional hazard model with the full set of interactions as the outcome model and the logit model as the propensity model. Appendix J provides details of the two models and of the cross-fitting procedure.
6.1 Optimal assignment policies
@@
Let $z_{ip}\in\left\{0,1\right\}$ be a variable indicating whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments maximizing the sum of treatment effects given by Equation 3:
@@
where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$. The first constraint ensures that the capacity limits are not violated, and the second one is that each applicant is assigned to one program. A policy that induces allocation $z_{ip}^{*}$ is what we call optimal policy. In Algorithm 1 in Appendix K, we provide further details of the estimation of optimal policy using training data, where we let $\hat{\pi}$ denote the estimate.
We estimate the optimal policy at three capacity levels: first, with the existing capacity (allowing 13% of all applicants into Mentoring and 15% into Challenges) – Optimal 1, second, increasing the capacity of Challenges to accommodate 50% of applicants – Optimal 2, and third, a scenario where each program can admit all applicants – Optimal 3. We compare these optimal policies against several alternatives: a policy accepting all applicants to Mentoring, a policy enrolling all applicants in Challenges (both these policies have increased capacities to accept all applicants), and the existing allocation policy (Status quo). We use the average treatment effect as the metric to compare policies. The results are presented in Figure 3.
Figure 3: Evaluation of counterfactual policies.
Note: Estimates of values of policies. The policy value is the mean probability of having a tech job across all participants and is measured by the mean of AIPW scores under allocation induced by each policy across all participants. Standard errors clustered at a participant level. Whiskers show confidence intervals. Optimal 1 has the existing capacity levels. Optimal 2 increases the capacity of Challenges to 50%, and Optimal 3 has no capacity limits.
First, we find that the optimal policy, with existing capacity levels, markedly improves the average treatment effect when compared to the Status quo, with an increase from 12 to 22 percentage points. Furthermore, this policy is also substantially more effective than policies that assign all applicants to Mentoring or to Challenges. This result follows from the high degree of treatment effect heterogeneity across both programs and shows that a targeting policy can substantially augment Dare IT’s effectiveness. Second, we find that the policy expanding the number of Challenges spots to 50% of applicants (Optimal 2) increases the average treatment effect to 17 percentage points and Optimal 3 to 15 percentage points. Such an increment highlights that the there are sizeable subpopulations amongst Dare IT applicants that have higher treatment effects. Moreover, we find only a small and statistically insignificant difference in policy value when contrasting Optimal 2 to the Optimal 3 policy.
In Table 5, we present the gains from optimal policy per assignment group. A program assignment group consists of all applicants that under a policy are assigned to this program, i.e., Mentoring assignment group are all applicants that are assigned to Mentoring. We are interested in comparing the benefits from the participation in each program of each of these groups under. We find that the group allocated to Mentoring substantially gains from their allocation, yielding an ATE of 0.199 (S.E. 0.017). This group would also substantially benefit from Challenges with the average treatment effect of 0.276 (S.E. 0.019). The difference between assigning them to Mentoring over Challenges is significant -0.077 (S.E. 0.021). However, the group assigned to Challenges benefits from this assignment even more. The average treatment effect of Challenges amounts to 0.244 (S.E. 0.017) compared to 0.109 (S.E. 0.012) of Mentoring. Lastly, the non-participating group has substantially lower benefits from assignment to Mentoring or to Challenges, average treatment effects of 0.101 (S.E. 0.006) and 0.073 (S.E. 0.003) respectively.
Table 5: Gains from the optimal assignment
Mentoring group Challenges group Nothing group
Share 0.15 0.50 0.35
@@
Optimal 0.199 (0.017) 0.244 (0.017)
Mentoring 0.199 (0.017) 0.109 (0.012) 0.101 (0.006)
Challenges 0.276 (0.019) 0.244 (0.017) 0.073 (0.003)
@@
Mentoring - Challenges -0.077 (0.021) -0.135 (0.025) 0.028 (0.006)
Note: Gains from the optimal assignment by assignment group from Optimal 1 - 13% of applicants get be admitted to Mentoring and 15% to Challenges. Groups are defined as all applicants assigned to the specific program. The columns describe the outcomes of groups under various allocations - the group membership is determined using Algorithm 1 and is constant across the comparisons here, the allocations change. The first row is the share of applicants assigned to a group. Rows two to four describe the average treatment effect of the group under various assignments. The 5th row is the average treatment effect under Mentoring less the average treatment effect under Challenges. Standard errors are next to each value in parentheses and are clustered at the applicant level.
When comparing the three policies based on optimal allocation, it’s evident that the current capacity levels limit Dare IT’s impact but that additional spots do not create more value beyond a certain capacity. To further unpack this, in Figure 4, we analyze the value of optimal policy under a sequence of capacity values and compare with random assignment to either Mentoring or Challenges. Purple dots in Figure 4, show the average probability of getting a job in technology across all applicants (including those not participating in Dare IT programs), for the total capacity of the two programs ranging from 0 to 100% (in all cases, the two programs have the same capacity, so for example the total capacity of 50% means that, up to 25% of applicants get be assigned to either program). We compare the outcome under these policies with policies that assign applicants randomly to Mentoring (in yellow) and to Challenges (in red).
Figure 4: The value of targeting at various capacity levels
Note: Estimates of mean outcomes across all applicants at various capacity levels of Dare IT programs and allocation mechanisms. The policy value is the predicted share of applicants getting a tech job. The share of applicants getting a spot refers to the total capacity in both programs. Challenges Random (in orange) and Mentoring Random (in red) assign applicants at random between the program and Out of Dare IT. Targeting policy (in purple) follow the optimal policy approach.
Our analysis indicates that policies randomly admitting applicants to either the Mentoring or Challenges programs lead to consistently lower policy values when compared to the targeted policy across all capacity levels. When capacities are low, targeted policies prioritize applicants with the highest benefits. A meaningful difference in the average probability of getting a job across all applicants quickly emerges between the targeted and random policies; consequently, already at the 25% capacity levels, there is an approximately 4 percentage points gap between targeted and random policies. Interestingly, even when capacity constraints are lifted to accommodate all applicants, the targeted strategy continues to lead to higher outcomes than the random policies based on one program. This discrepancy remains because there of the difference in benefits that applicants receive across the two programs, i.e., there are some applicants that benefit more from Challenges than they do from Mentoring.
6.2 Counterfactual selection policies for mentors

### v4

@@ 6 Counterfactual Program Assignment Policies
The number of spots in each program is limited. The availability of mentors restricts the capacity of Mentoring. In Challenges, we limited the available spots to 300 to accommodate potential operational complexities and logistical problems. Considering the substantial treatment effects of both programs, relaxing these capacity constraints will increase Dare IT’s impact. However, capacity constraints are likely to remain a long-term issue for the Mentoring program and may persist, albeit to a lesser extent, for Challenges. Consequently, Dare IT will continue to face the question of which applicants to accept for each program. Given the heterogeneous treatment effects, admission policies that favor applicants with high predicted benefits could enhance the impact of Dare IT programs. This section aims to quantify these potential gains.
@@
We perform three analyses: (i.) we estimate gains from the introduction of a program admission policy that, given capacity levels in each program, assigns applicants to programs, maximizing the sum of predicted treatment effects of programs’ participants; (ii.) we evaluate the gains from relaxing capacity constraints with the targeting policy in place; (iii.) we compare the overall impact of Mentoring when mentors select the most promising candidates to a case where they select to following alternative strategies.
An assignment policy is a mapping from the characteristics of the applicants $X\in\mathcal{X}$ and the capacity levels of the programs $Q=(Q^{M},Q^{C})\in\mathcal{Q}$ to assignments: $\pi:\mathcal{X}\times\mathcal{Q}\rightarrow\mathcal{P}$, where $\mathcal{P}\equiv\left\{M,C,O\right\}$ and $M$ stands for Mentoring, $C$ for Challenges, and $O$ for Out of Dare IT. $\mathcal{Q}$ is the grid of possible capacity levels under consideration.
Policy estimation and evaluation are performed separately in train and test sets. Across all analyses, we obtain train and test sets by randomly splitting data at the applicant level with equal probability ($\mathcal{I}_{train}$ and $\mathcal{I}_{test}$ are resulting sets of applicants in the train and test data sets, respectively). The objective of policy estimation is to obtain a mapping from applicant characteristics and programs’ capacity levels to program assignments. We use the train set to obtain the estimate of the policy $\pi$, which we denote $\hat{\pi}$.
@@
Next, to evaluate a policy, we proceed in two steps: first, we apply the estimated mapping $\hat{\pi}$ to the test set, where for each $i$ in the test set and $Q\in\mathcal{Q}$, we obtain the counterfactual assignment of the applicants to programs $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. Second, we construct the AIPW estimate of the “policy value,” that is, the outcome value when treatments are assigned according to $\hat{\pi}(X_{i},Q)\in\mathcal{P}$. We do that using cross-fitting at the applicant level in the test sample, i.e., we estimate new outcome and propensity models and obtain AIPW estimates of the policy values in held-out folds; specifically, we use 100 repetitions with 5 folds and consider the median outcome across repetitions in (chernozhukov2017double).
Throughout this section, we combine the two experimental data sets with data on non-selected applicants to Mentoring and late applicants to Challenges. Thus, we have 997 applicants who did not participate in any program within our sample, resulting in a total sample of 1314 applicants and 19,000 applicant-month observations. For both policy estimation and evaluation, we use survival forests to estimate outcome models and regression forests for the propensity models. The propensity model adjusts for differences between participants and non-participants determined by selection to Mentoring and differences in the populations of applicants to Mentoring and to Challenges. Appendix H provides details of the models used in policy estimation. Appendix I, presents details of the models used in the policy evaluation, the cross-fitting procedure, and validates the approach by comparing the off-policy estimates with the randomized controlled trial estimates. Our procedure recovers the experimental estimates.
Outcome measure.
In all the preceding analyses, we measured treatment effects as changes in the probability of securing technology employment. For policy evaluation, we adopt an alternative metric that captures additional dimensions of program impact: the restricted mean survival time (RMST), which measures the average time until technology job acquisition over our observation period. The RMST metric offers two advantages for comparing policies. First, it incorporates both whether participants find jobs and how quickly they do so, a policy that helps participants find jobs faster, even if it doesn’t change the ultimate employment probability, generates economic value. Second, RMST estimates aggregate more naturally across heterogeneous populations and policy scenarios, facilitating comparisons of the assignment rules. We report RMST gains in months: negative values indicate participants find technology employment sooner relative to non-participants. An RMST gain of -3 months means treated participants find technology jobs an average of three months earlier. We use survival forests to estimate average treatment effects (athey2019generalized; cui2023estimating).
6.1 Targeted Assignment Policies
@@
Let $z_{ip}\in\left\{0,1\right\}$ be a variable that indicates whether applicant $i$ is assigned to program $p\in P$, and $z_{ip}^{*}$ be the arguments that maximize the sum of treatment effects given by Equation 2:
@@
where $\tau_{i}^{p}=\mathbb{E}\left[Y(p)-Y(O)\right]$. The first constraint ensures that capacity limits are not violated, and the second is that each applicant is assigned to one allocation. A policy that induces allocation $z_{ip}^{*}$ is what we call the targeted policy. $\tau_{i}^{p}$’s are estimated using the AIPW estimator trained on the train set. In Algorithm 1 in Appendix H, we provide further details of the estimation of the targeted policy.
The outcome of interest, policy value, is the participants’ average treatment effect given program allocation induced by the policy $\pi$. Specifically, we use the AIPW estimator to estimate the average difference in outcomes between being selected to participate based on the assignment policy and not participating in any Dare IT program. This emphasis on impact, rather than average outcomes of participants, aligns with Dare IT’s mission to create the greatest benefit for their participants.
6.2 Value of Targeted Policies
Given limited program capacity, which applicants should be assigned to which program to maximize overall impact? We estimate the targeted policy at three capacity levels: the 15/15/70 policy approximates existing capacity (assigning 15% of all applicants to Mentoring and 15% to Challenges); the 33/33/34 policy increases capacity to 33% for each program; and the 50/50/0 policy allows each program to admit 50% of applicants, ensuring a program slot for every applicant. We compare these policies against random allocation as a benchmark (random policy allocates all applicants at random to either Mentoring or Challenges). In Section 6.3, we examine how mentor-based selection compares to the algorithm-driven policy in greater detail. The policies vary in both capacity levels and whether participants are systematically targeted.
To understand which applicant types are assigned to each program under data-driven targeting, Figure 3 presents standardized differences in baseline characteristics between optimally assigned individuals and those not assigned to treatment. We implement the 15/15/70 targeted allocation policy and compare covariate distributions of each assigned group to the not-assigned group. Circles represent comparisons for Mentoring; triangles represent Challenges. Dashed vertical lines mark the $\pm$0.1 threshold for imbalance.
Figure 3: Characteristics of Optimally Assigned Groups
Notes: Standardized differences in baseline characteristics between optimally assigned individuals and those not assigned under the 15/15/70 allocation policy. Circles show Mentoring vs Not Assigned; triangles show Challenges vs Not Assigned. The solid vertical line at zero indicates perfect balance; dashed lines mark the $\pm$0.1 threshold. Standardized differences are calculated as the difference in means divided by the pooled standard deviation. Positive values indicate the optimally assigned group has higher values on that characteristic.
The targeted allocation exhibits clear differentiation in targeting patterns across programs. Applicants selected for Mentoring are more likely to hold STEM degrees, live outside Warsaw, have Master’s degrees, and be over 30 years old. Notably, although the algorithm does not use information about prior mentor selection, previously mentor-selected applicants are disproportionately targeted by the algorithm—suggesting the model captures similar signals to those used by mentors. For Challenges, the algorithm shows more balanced selection across characteristics. Assigned applicants are moderately more likely to have been previously selected by mentors but are less likely to hold STEM degrees, live in small towns, or have extensive professional experience compared to those not selected.
Main result.
Table 6 shows gains from data-driven targeting, with effect sizes varying by program and capacity constraint. Under the most selective policy (15/15/70), the overall treatment effect reaches 4.34-4.34 months, which corresponds to an 86% improvement over random allocation’s 2.33-2.33 months. These gains exhibit a strong capacity-selectivity tradeoff: as capacity expands from 15% to 50%, per-person treatment effects decline (from 4.34-4.34 to 2.86-2.86 months overall), though even universal coverage (50/50/0) achieves effects 23% larger than random allocation. Thus, we find that there are gains from targeting across programs.
The targeting gains differ across programs. Mentoring exhibits the largest effects under selective policies (5.55-5.55 months at 15% capacity versus 2.31-2.31 under random), while Challenges shows more modest differentiation (3.13-3.13 versus 2.35-2.35). Notably, even at 50% capacity where the algorithm must accept marginal applicants with lower predicted treatment effects, Mentoring maintains treatment effects 54% larger than random allocation, demonstrating that observable characteristics contain substantial predictive power for treatment response beyond what random assignment captures.
Table 6: Targeted Allocation Policies: Treatment Effects by Program
Capacity Allocation Mentoring Challenges Overall
(Ment. / Chal. / None) Effect Effect Effect
Panel A: Benchmark
Random Allocation -2.31 (0.022) -2.35 (0.021) -2.33 (0.022)
Panel B: Targeted Allocation Policies
15% / 15% / 70% -5.55 (0.086) -3.13 (0.069) -4.34 (0.055)
33% / 33% / 34% -4.33 (0.052) -2.68 (0.041) -3.50 (0.033)
50% / 50% / 0% -3.55 (0.039) -2.17 (0.027) -2.86 (0.024)
Notes: Treatment effects measured as change in months to employment (negative values = faster job finding). Standard errors in parentheses. Each row shows a capacity-constrained allocation policy specifying the share of applicants assigned to Mentoring, Challenges, or neither program. Random Allocation assigns each program to a random 50% of applicants.
Table 7 examines treatment effects within each assignment group under the 15/15/70 targeted policy. An assignment group consists of all applicants allocated to a specific program (Mentoring, Challenges, or neither) under this policy. For each group, we estimate counterfactual treatment effects under alternative program assignments to assess whether the targeted allocation correctly matches individuals to their highest-benefit program.
The results reveal effective targeting. The Mentoring assignment group receives large benefits from their targeted allocation (5.55-5.55 months, SE = 0.086), substantially exceeding what they would gain from Challenges (3.12-3.12 months, SE = 0.062). The difference of 2.43 months (SE = 0.106) is statistically significant. The Challenges assignment group also benefits most from their assigned program (3.13-3.13 months, SE = 0.069) compared to Mentoring (1.87-1.87 months, SE = 0.050); a difference of 1.26 months (SE = 0.085). This pattern demonstrates benefits from targeting across both programs rather than only selecting to a program.
The Out of Dare IT group (70% of applicants receiving no program assignment) exhibits substantially smaller treatment effects under either counterfactual assignment: 1.72-1.72 months (SE = 0.024) for Mentoring and 2.02-2.02 months (SE = 0.022) for Challenges. While this group would experience gains from program participation, their predicted effects are smaller than those optimally assigned to treatment.
Table 7: Gains From Assignment Policy per Assignment Group
Mentoring group Challenges group Out of Dare IT group
Share 0.15 0.15 0.70
@@ 6.2 Value of Targeted Policies
Targeted -5.552 (0.086) -3.132 (0.069)
Mentoring -5.552 (0.086) -1.872 (0.050) -1.720 (0.024)
Challenges -3.122 (0.062) -3.132 (0.069) -2.017 (0.022)
@@
Mentoring - Challenges -2.431 (0.106) 1.259 (0.085) 0.297 (0.032)
Note: ATEs are cross-fitted AIPW estimates from causal survival forests (RMST changes up to the evaluation horizon). Negative values indicate earlier job finding (months sooner) relative to out of Dare IT programs. Standard errors in parentheses are GRF mean-of-variance plug-in SEs aggregated within arm.
Policy value across capacity levels.
When comparing the average treatment effects under targeted assignments and the random assignment, we find that there are subsets of applicants who benefit substantially from targeting. However, when the capacity of the two programs increases, the additional participants who get in have lower treatment effects, as suggested by the comparison of targeted policies in Table 6. Thus, there are diminishing returns to targeting additional slots. To identify the level of capacity that realizes most of the available gains from targeting, we estimate the outcomes from targeted assignment under a sequence of capacity values and compare the resulting average probability of getting a job in technology with this outcome under random assignments to Mentoring and to Challenges.
Figure 4: Policy Value Across Capacity Levels: Targeted vs Random Allocation
Notes: Average treatment effects under targeted and random allocation policies as capacity varies from 5% to 50% per program. The solid curve shows targeted allocation that assigns individuals with highest predicted treatment effects to each program; the dashed line shows random assignment of 50% to each program. Treatment effects measured as change in months to employment (negative = faster job finding). Shaded region indicates 95% confidence interval for targeted policy. Policy value evaluated using cross-fitted AIPW estimates on test set. At each capacity level, targeted allocation maximizes expected policy value by solving the assignment problem in Equation 2.
Figure 4 presents the capacity-selectivity tradeoff by plotting average treatment effects as program capacity varies from 5% to 50% per program. The solid curve shows targeted allocation; the dashed horizontal line represents random assignment. Several patterns emerge. First, targeted allocation consistently outperforms random allocation across all capacity levels, with gains largest when capacity is most constrained. At 15% capacity per program, the targeted policy achieves treatment effects of 4.3-4.3 months compared to random allocation’s 2.3-2.3 months, a 2-month improvement. Second, the targeted policy exhibits sharply diminishing returns as capacity expands: per-person effects decline from approximately 4.5-4.5 months at 5% capacity to 2.9-2.9 months at 50% capacity. This reflects the algorithm’s sequential assignment of highest-benefit individuals first, with marginal participants exhibiting progressively smaller treatment effects. Third, even at universal coverage (50% per program, accommodating all applicants), targeted allocation maintains a 0.6-month advantage over random assignment—demonstrating that complementarity between individual characteristics and program type generates value beyond pure capacity expansion.
6.3 Counterfactual Mentors’ Selection Policies
Applicants selected by mentors and randomly assigned to the control group had a significantly higher likelihood of securing a tech job compared to those who were not selected (see Figure 2). This suggests that mentors identify promising applicants who are likely to secure a tech job regardless of their participation in Mentoring. Table 8 compares characteristics of selected and non-selected applicants (panel A), and selected by different groups of mentors (panels B and C). Baseline is the share of control group applicants that found a job.
Table 8: Differences in Applicant Characteristics: Selection and Mentor Experience
Baseline Small town Warsaw Over 30 Prof. exp. Social science STEM Applied before Mother High skill Family/friends in IT
Panel A: Selected vs. Not Selected
Selected 0.456 0.245 0.268 0.487 7.097 0.134 0.510
(0.041) (0.025) (0.026) (0.029) (0.284) (0.020) (0.029) (–) (–) (–) (–)
Not 0.286 0.476 0.282 0.147 7.489 0.207 0.182
(0.021) (0.023) (0.021) (0.016) (0.288) (0.019) (0.018) (–) (–) (–) (–)
Panel B: Mentor Managerial Experience
High 0.378 0.229 0.281 0.588 7.595 0.137 0.497 0.261 0.288 0.614 0.699
(0.057) (0.034) (0.036) (0.040) (0.417) (0.028) (0.041) (0.036) (0.037) (0.039) (0.037)
Low 0.549 0.266 0.245 0.378 6.503 0.133 0.531 0.210 0.231 0.497 0.699
(0.059) (0.037) (0.036) (0.041) (0.384) (0.028) (0.042) (0.034) (0.035) (0.042) (0.038)
Panel C: Mentor Total Experience (above vs. below median)
Above 0.390 0.239 0.299 0.590 8.197 0.171 0.564 0.282 0.308 0.573 0.718
(0.064) (0.040) (0.043) (0.046) (0.442) (0.035) (0.046) (0.042) (0.043) (0.046) (0.042)
Below 0.512 0.251 0.240 0.419 6.330 0.112 0.480 0.207 0.229 0.547 0.687
(0.054) (0.033) (0.032) (0.037) (0.364) (0.024) (0.037) (0.030) (0.031) (0.037) (0.035)
Notes: Means shown with standard errors in parentheses. Baseline refers to the average outcome when not participating in Mentoring. Panel A compares applicants who were selected vs. not selected. Panel B compares applicants selected by mentors with vs. without managerial experience. Panel C compares applicants selected by mentors above vs. below the median years of experience.
Panel A of Table 8 reveals systematic differences between applicants selected by mentors and those not selected. Selected applicants exhibit substantially stronger baseline employment prospects (0.456 vs 0.286), despite having similar years of professional experience (7.1 vs 7.5 years). Selected applicants are more likely to be over 30 years old (0.487 vs 0.147) and to hold STEM degrees (0.510 vs 0.182), while being considerably less likely to come from small towns (0.245 vs 0.476). They are also less likely to have social science backgrounds (0.134 vs 0.207).
Panel B compares mentors with versus without managerial experience. Mentors with managerial experience select applicants with weaker baseline employment prospects (0.378 vs 0.549) but who are substantially older (0.588 vs 0.378) and have more professional experience (7.6 vs 6.5 years). Panel C shows that mentors with above-median total career experience exhibit similar but more pronounced patterns. They select applicants with weaker baselines (0.390 vs 0.512) who are older (0.590 vs 0.419), have substantially more professional experience (8.2 vs 6.3 years), and are more likely to hold a STEM degree (0.564 vs 0.480). The consistency across Panels B and C suggests a common selection philosophy among more experienced mentors: they look beyond current employability to identify "high-potential, high-barrier". Less experienced mentors appear to select more conventionally, favoring applicants already demonstrating stronger baseline employability.